OpenStudy (jiteshmeghwal9):

If the ratio between the roots of the equation , \(lx^2+nx+n=0\) is P:Q then the value of \(\sqrt{p/q}+\sqrt{q/p}+\sqrt{n/l}\) is :-

OpenStudy (anonymous):

Let the roots be a and b...

OpenStudy (jiteshmeghwal9):

k @Yahoo!

OpenStudy (anonymous):

\[\frac{x_1}{x_2}=\frac{p}{q}\]\[(\sqrt{\frac{p}{q}}+\sqrt{\frac{q}{p}})^2=(\sqrt{\frac{x_1}{x_2}}+\sqrt{\frac{x_2}{x_1}})^2=\frac{x_1}{x_2}+\frac{x_2}{x_1}+2=\frac{(x_1+x_2)^2}{x_1x_2}=\frac{n}{l}\]does that help?